# Phase Selection and Analysis for Multi-frequency Multi-user RIS Systems Employing Subsurfaces in Correlated Ricean and Rayleigh Environments

### v1

@@ Abstract
We analyse the performance of a reconfigurable intelligent surface (RIS) aided system where the RIS is divided into subsurfaces. Each subsurface is designed specifically for one user, who is served on their own frequency band. The other subsurfaces (those not designed for this user) provide additional uncontrolled scattering. We derive the exact closed-form expression for the mean signal-to-noise ratio (SNR) of the subsurface design (SD) when all channels experience correlated Ricean fading. We simplify this to find the mean SNR for line-of-sight (LoS) channels and channels experiencing correlated Rayleigh fading. An iterative SD (ISD) process is proposed, where subsurfaces are designed sequentially, and the phases that are already set are used to enhance the design of the remaining subsurfaces. This is extended to a converged ISD (CISD), where the ISD process is repeated multiple times until the SNR increases by less than a specified tolerance. The ISD and CISD both provide a performance improvement over SD, which increases as the number of RIS elements (N) increases. The SD is significantly simpler than the lowest complexity multi-user (MU) method we know of, and despite each user having less bandwidth, the SD outperforms the existing method in some key scenarios. The SD is more robust to strongly LoS channels and clustered users, as it does not rely on spatial multiplexing like other MU methods. Combined with the complexity reduction, this makes the SD an attractive phase selection method.
Index Terms:
Reconfigurable Intelligent Surface (RIS), array signal processing, correlated Ricean and Rayleigh fading
@@ I Introduction
Reconfigurable Intelligent Surfaces (RIS) are the focus of considerable research interest for next generation mobile communications, due to their ability to manipulate the wireless channel with minimal power consumption. RIS can enhance several aspects of system perfomance, including signal-to-noise-ratio (SNR), rank deficiency and blockage avoidance [1]. There is now a significant body of work on RIS, including design, testbeds, and performance analysis [2]. Standards development is also underway, with ETSI releasing its first report on RIS in 2023 [3].
Typical RIS implementations involve large numbers of elements, $N$, in an array, where each element can be set to control the phase of the reflected signal. A key challenge is selecting the RIS element reflection coefficients to maximise the total SNR at the target receiver. An analytically determined optimal RIS design does not appear feasible for a multi user (MU) system. Iterative schemes have been designed [4, 5, 6, 7], but have high computational complexity. In order to keep power consumption low, limited processing can be done on the RIS itself. In addition, these methods demand significant amounts of channel state information (CSI) and necessitate complex receiver processing. Typically, the matrix operations required are similar to those for minimum mean squared error (MMSE) receivers. Therefore, while they may perform well, these methods would be challenging to implement.
Complexity reduction research in RIS design includes using statistical CSI, such as in the codebook designs in [8], and the study of the trade off between energy efficiency and spectral efficiency in [9]. In terms of using channel estimates, lower complexity has been achieved for both a double RIS system [10] and a discrete phase shift system [11]. While both considerably reduce overall computation, they target high performance. This necessitates iterative optimisation and repeated matrix operations (inversion or singular value decomposition (SVD)). In this paper, we take a different direction to these works, targeting useful (but not near-optimal) performance improvements for the RIS while keeping the RIS design and receiver processing remarkably simple.
@@
In [12], we proposed a low complexity subsurface design (SD) where a RIS of $N$ elements is divided into $K$ subsurfaces with $N_{1},N_{2},\dots,N_{K}$ elements, respectively, where $K$ is the number of users. Each user operates on a different frequency, so that matched filtering (MF) can be used at the receiver, and one subsurface is designed per user. The elements designed for a specific user are selected using the optimal single user (SU) phase selection method proposed in [13]. The elements designed for other users provide uncontrolled scattering for this user. This may assist the user, particularly in higher frequency systems where there is little natural uncontrolled scattering in the environment. As elements are only being designed for one of $K$ users, channel estimation is reduced by a factor of $K$. The use of MF ensures that the receiver processing is as simple as possible. Note that simple beamforming is more practical at higher frequencies and thus SU systems remain of strong interest in 5G [14].
@@
In our preliminary work in [12], we only considered the performance of the SD for a line-of-sight (LoS) RIS-base station (BS) channel and correlated Rayleigh user equipment (UE)-BS and UE-RIS channels. In reality, it is likely that all channels will have both LoS and non-line-of-sight (NLoS) components. Therefore, we now significantly extend our work by investigating the performance of this simple and practical method for more general and realistic correlated Ricean fading channels. This versatile analytical tool can be simplified to more specific combinations of LoS and NLoS channels as desired. We also extend the subsurface design to a low complexity iterative subsurface design (ISD). The SD assumes that each subsurface is designed independently with no knowledge of phases set for other users. The ISD sets the subsurfaces sequentially from the weakest to strongest user, utilising the knowledge of all phases set for other users to improve the design. Thus, users with the largest impact on total SNR benefit from the most prior knowledge. The ISD sees this process performed once, but it can be performed repeatedly until it converges and the total SNR stops increasing. This is referred to as the converged ISD (CISD). The ISD and CISD offer performance improvements while still benefiting from relatively low computational complexity and a simple MF receiver.
@@
Using analysis and simulations, we both verify the accuracy of our derivation and investigate the performance of the SD and ISD. The mean SNR of both methods is compared for a range of RIS dimensions. The mean sum rate supported by the SD is compared to a lowerbound benchmark equivalent to no channel knowledge (e.g. randomly selected RIS phases) and the MU method in [4]. This was selected for comparison as it is the lowest complexity existing full MU design that we are aware of, and while other methods may offer better performance, it is highly probable that they will require considerably higher system complexity. The impact of a range of system parameters is considered including Ricean K-factor, correlation and array dimensions.
@@
Notation: Upper and lower boldface letters represent matrices and vectors, respectively. We define vector $\mathbf{v}=[\mathbf{v}_{1}^{T},\dots,\mathbf{v}_{K}^{T}]^{T}$, where $\mathbf{v}_{k}$ is the $k$-th block of vector $\mathbf{v}$ and $\mathbf{v}_{k,i}$ is the $i$-th element of the $k$-th block. $\mathbf{M}$ is a matrix comprising of submatrices, or blocks. $\mathbf{M}_{r,s,i,j}$ is the $(i,j)$-th element of the $r$-th block across and $s$-th block down. Matrices are indexed this way except where explicitly defined otherwise in the text. $\mathbf{M}^{(k)}$ denotes a matrix relating to user $k$. $\mathbb{C}$ is the set of complex numbers. $\Re[\cdot]$ is the Real operator. $\mathcal{CN}(\boldsymbol{\mu},\mathbf{Q})$ represents a complex Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance matrix $\mathbf{Q}$. $\mathbb{E}[\cdot]$ represents statistical expectation. $(\cdot)^{T}$, $(\cdot)^{*}$ and $(\cdot)^{\dagger}$ represent the transpose, conjugate and Hermitian transpose operators, respectively. The angle of a vector, $\mathbf{x}$, of length $N$ is denoted as $\angle\mathbf{x}=[\angle{x}_{1},\dots,\angle{x}_{N}]^{T}$ and the exponent as $e^{\mathbf{x}}=[e^{{x}_{1}},\dots,e^{{x}_{N}}]^{T}$. |||||$\cdot$||| | ⋅ | | is the Euclidean norm. ${}_{1}F_{1}(a,b;z)$ is the confluent hypergeometric function and ${}_{2}F_{1}(a,b,c;z)$ is the Gaussian hypergeometric function. $L_{\nu}(\cdot)$ denotes a Laguerre function of non-integer degree $\nu$. $\Gamma(\cdot)$ is the complete gamma function. $f^{*}(r,\theta)$ is the probability density function (PDF) of a complex random variable $X=r\,\mathrm{e}^{j\theta}$, and $f^{*}(r_{1},r_{2},\theta_{1},\theta_{2})$ is the joint PDF of complex random variables $X_{1}=r_{1}\mathrm{e}^{j\theta_{1}}$ and $X_{2}=r_{2}\mathrm{e}^{j\theta_{2}}$. $J_{n}(\cdot)$ is the $n$-th order Bessel function of the first kind and $I_{n}(\cdot)$ is the $n$-th order modified Bessel function of the first kind.
@@ II System Model
We consider an uplink RIS aided system, as shown in Fig. 1, consisting of a RIS panel with N elements and a BS with M antennas. K single-antenna users are located in the vicinity of the BS and RIS. A RIS link (UE - RIS - BS) and a direct link (UE - BS) connect the UE to the BS. The system bandwidth, $B$, is split into $K$ bands of $\frac{B}{K}$ Hz, with one user per band. The phase response of the RIS, and therefore the phase shift in the reflected signal, is fixed across the frequency bands. The RIS elements are grouped into “subsurfaces” and each subsurface is designed for a different user.
Figure 1: System model showing uplink channels for UE 1 in band 1.
For user $k$ in band $k$, let $\mathbf{h}^{(k)}_{\mathrm{d}}\!\in\mathbb{C}^{M\times 1}$, $\mathbf{h}^{(k)}_{\mathrm{ur}}\in\mathbb{C}^{N\times 1}$ and $\mathbf{H}^{(k)}_{\mathrm{rb}}\in\mathbb{C}^{M\times N}$ be the direct, UE-RIS and RIS-BS channels, respectively. $\mathbf{\Phi}\in\mathbb{C}^{N\times N}$ is a diagonal matrix of reflection coefficients for the RIS which can be given in block diagonal form, where the $k$-th block is designed to enhance the channel for user $k$. $\mathbf{\Phi}=\mathrm{diag}\{\mathbf{\Phi}_{1},\dots,\mathbf{\Phi}_{K}\}$, where $\mathbf{\Phi}_{k}=\mathrm{diag}(e^{j\phi_{k,1}},\dots,e^{j\phi_{k,N_{k}}})$. $N_{k}$ is the number of RIS elements chosen to support user $k$ and $\sum^{K}_{k=1}N_{k}=N$. The received signal in band $k$ at the BS is
@@
where $s_{k}$ is the signal being sent from user $k$, $\mathbb{E}[|s_{k}|^{2}]=E_{s}$, and $\mathbf{n}_{k}\sim\mathbb{C}\mathcal{N}(0,\sigma^{2}\mathbf{I})$ is additive white Gaussian noise. Using the block diagonal form of $\mathbf{\Phi}$, (1) becomes
@@ II-A Channel Model
For this work, we consider correlated Ricean channels consisting of a rank-1 LoS and correlated Rayleigh component for all links. The LoS components are comprised of steering vectors and the scattered components adopt the Kronecker correlation model. Therefore, the channels for user $k$ are
@@
and where $\beta^{(k)}_{\mathrm{d}}$, $\beta^{(k)}_{\mathrm{rb}}$ and $\beta^{(k)}_{\mathrm{ur}}$ are the channel gains, $\kappa^{(k)}_{\mathrm{d}}$, $\kappa^{(k)}_{\mathrm{rb}}$ and $\kappa^{(k)}_{\mathrm{ur}}$ are the Ricean K-factors, $\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{LoS})}$, $\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{LoS})}$ and $\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{LoS})}$ are the LoS channel components for the $k$th user and $\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{SC})}$, $\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{SC})}$ and $\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{SC})}$ are the scattered channel (SC) components for the $k$th user. Here,
@@
This work focuses on strongly LoS $\mathbf{H}_{\mathrm{rb}}$ channels, where one ray is dominant but a scattered component is also present. This is a common and reasonable assumption for RIS systems, as the RIS can be located to exploit the LoS channel as seen in [15, 16, 17].

### v2

@@ Abstract
Phase selection design for reconfigurable intelligent surfaces (RISs) is a significant research challenge, as a closed-form optimal solution for a multi-user (MU) system is believed to be intractable. While existing methods achieve strong near-optimal performance, they typically entail high computational complexity. In this work, we take a different approach and propose a practical method that achieves competitive performance while substantially reducing computational complexity. To do so, we consider a RIS divided into subsurfaces. Each subsurface is designed specifically for one user, who is served on their own frequency band. The other subsurfaces (those not designed for this user) provide additional uncontrolled scattering. We derive the exact closed-form expression for the mean signal-to-noise ratio (SNR) for the proposed subsurface design (SD) when all channels experience correlated Ricean fading. We simplify this to find the mean SNR for line-of-sight (LoS) channels and channels experiencing correlated Rayleigh fading. An iterative SD (ISD) process is proposed, where subsurfaces are designed sequentially, and the phases that are already set are used to enhance the design of the remaining subsurfaces. This is extended to a converged ISD (CISD), where the ISD process is repeated multiple times until the SNR increases by less than a specified tolerance. The ISD and CISD both provide a performance improvement over SD, which increases as the number of RIS elements increases. The SD is significantly simpler than the lowest complexity MU method we know of, and despite each user having less bandwidth, the SD outperforms the existing method in some key scenarios. The SD is more robust to strongly LoS channels and clustered users, as it does not rely on spatial multiplexing like other MU methods. Combined with the complexity reduction, this makes the SD an attractive phase selection method.
@@ I Introduction
Reconfigurable intelligent surfaces (RISs) are the focus of considerable research interest for next generation mobile communications, due to their ability to manipulate the wireless channel. RIS can enhance several aspects of system performance, including signal-to-noise ratio (SNR), rank deficiency, and blockage avoidance [1]. A significant body of work on RIS now exists, including design, testbeds, and performance analysis [2]. Standards development is underway, with ETSI releasing its first report on RIS in 2023 [3].
Typical RIS implementations involve large numbers of elements in an array, where each element can be set to control the phase of the reflected signal. A key challenge is selecting the RIS element reflection coefficients to maximize the total SNR at the target receiver. Analytically optimal closed-form RIS designs for multi-user (MU) systems are not tractable, and many existing methods heavily rely on iterative optimization [4, 5, 6, 7, 8, 9, 10]. While these offer excellent performance, they involve significant computational overhead. This becomes particularly challenging when the RIS is passive and lacks onboard processing. Moreover, passive RIS cannot transmit pilots or process feedback, complicating channel estimation, particularly as the number of users increases. Therefore, reducing both the computational and channel estimation requirements is essential for the practical deployment of RIS-assisted MU systems.
I-A Review of Related Literature
The design of RIS phase selection methods has primarily focused on either iterative optimization methods, or the use of artificial intelligence (AI) and machine learning (ML) approaches, as analytically optimal closed-form RIS designs for MU systems are not tractable.
The foundational work in [4] presents a sum rate maximization method for MU RIS systems, where the RIS phase shift matrix and the base station (BS) precoding matrix are alternately optimized while considering non-uniform RIS reflection amplitudes. A similar problem was solved in [5] with imperfect channel state information (CSI).
In [6], the authors proposed a weighted downlink (DL)uplink (UL) sum rate optimization framework in which the DL and UL precoding matrices and the RIS phase shift matrix are alternately optimized. For a single-BS system, [7] proposes multiple joint optimization strategies over the RIS phases and BS beamforming vectors to maximize the SNR, and for a dual-BS system, the authors propose a joint optimization approach over the RIS phases and BS transmit powers that maximizes the mean signal-to-interference-plus-noise ratio (SINR). In [8], the transmit power is minimized by alternately optimizing the RIS phase shifts (in this case selected from a discrete set) and the BS beamforming vector under imperfect CSI.
The works in [9, 10, 11, 12] focus on lower complexity iterative methods. In [9], a method is proposed that selects the RIS phases to increase the sum rate by reducing the total mean-squared error (MSE) of the receiver rather, and in [10], the precoder, combiner and RIS phases are jointly optimized to minimize the MSE. The work in [11] proposes a low complexity approach to RIS channel estimation and the optimization of discrete RIS phases to maximize the sum rate, and in [12], a low complexity joint optimization method is detailed for a system involving two RISs, where the RIS phases and BS beamforming are jointly optimized to maximize sum rate.
The application of AI/ML to the RIS phase selection problem has allowed optimal phases to be found for more complex and larger systems. The work in [13] trains and tests a fully connected neural network to find the RIS phases that minimize the bit error rate (BER) and maximize the sum rate even with imperfect CSI. The authors of [14] use deep reinforcement learning to jointly optimize the RIS phases and power allocation coefficients to solve a multi-objective optimization problem, where power consumption is minimized and data rate is maximized. In [15], unsupervised learning is used to jointly optimize the BS beamforming vector and the RIS phases to maximize the SNR in a large-scale RIS system with thousands of elements.
All of the methods discussed above are designed to achieve optimal or near-optimal performance, and they deliver excellent results. However, even the low-complexity approaches within these [9, 10, 11, 12] rely on iterative procedures and non-trivial matrix operations and optimization steps. As a result, their computational burden can still be significant.
I-B Contributions
In this paper, we take a different direction to the works discussed in the previous section, and target useful performance gains with an extremely low computational complexity RIS phase selection method in a practical orthogonal frequency-division multiplexing (OFDM)-based system. While large-scale systems with multiple RISs are an important topic for research, the availability of exact analytical results for RIS systems remains limited. Thus, beginning with a simpler case allows general system insights to be gained through exact expressions and can provide a foundation for analyzing more complex systems in the future.
@@ I-B Contributions
In [16], we proposed a low-complexity subsurface design (SD) where a RIS is divided into subsurfaces and one subsurface is designed to serve one user equipment (UE). A MU system employing OFDM is considered. OFDM systems are ubiquitous in 4G and 5G systems, and are expected to be so in 6G systems also [17]. Each user operates on a different frequency, which enables simple matched filtering (MF) to be used at the receiver and removes inter-user interference (IUI). The elements designed for a specific UE are selected using the optimal single-user (SU) phase selection method proposed in [18], and are optimal for that UE due to the lack of IUI. The elements designed for other UEs provide uncontrolled scattering for this UE. This may assist the UE, particularly when there is little natural uncontrolled scattering in the environment. As the elements are designed for a single UE, only the channels from that UE through one RIS subsurface need to be estimated. This leads to a reduction in the channel estimation requirement proportional to the number of UEs compared with designs where all UEs operate in the same frequency band. The use of MF ensures that the receiver processing is as simple as possible. The SD is not intended to compete with the most advanced iterative methods, but rather makes an important contribution to a critical gap in RIS research, namely simple low-complexity MU phase selection schemes that are viable for real-world implementation.
Motivated by the identified research gaps, we significantly extend the preliminary work in [16] by broadening the analysis conducted to a much more general scenario, and proposing multiple extensions to the SD that offer further performance gains. The main contributions of this work are as follows:
@@
  • In our preliminary work in [16], we only considered the performance of the SD for a line-of-sight (LoS) RIS-BS channel and correlated Rayleigh UE-BS and UE-RIS channels. In reality, it is likely that all channels will have both LoS and non-line-of-sight (NLoS) components. Therefore, we now significantly extend our work by investigating the performance of this simple and practical method for more general and realistic correlated Ricean fading channels. This versatile analytical tool can be simplified to more specific combinations of LoS and NLoS channels as desired.
  • We also extend the subsurface design to a low-complexity iterative subsurface design (ISD). The SD assumes that each subsurface is designed independently with no knowledge of phases set for other users. The ISD sets the subsurfaces sequentially, utilizing the knowledge of all phases set for other users to improve the design. The ISD performs this process once, but it can be performed repeatedly until convergence and the total SNR stops increasing. This is referred to as the converged ISD (CISD). The ISD and CISD offer performance improvements while still benefiting from low computational complexity and a simple MF receiver.
@@
  • Using analysis and simulations, we investigate the performance of the SD, ISD and CISD. The mean SNR of each method is compared for a range of RIS dimensions. The mean sum rate supported by the SD is compared to a lower-bound benchmark equivalent to no channel knowledge (e.g. randomly selected RIS phases) and the MU method in [9]. This was selected for comparison as it is the lowest complexity existing full MU design that we are aware of, and while other methods may offer better performance, it is highly probable that they will require considerably higher system complexity. We thoroughly investigate the difference between the SD and MU method for a wide range of system parameters, including RIS size, correlation between BS and RIS elements, Ricean K-factor, UE location and UE clustering, and highlight important scenarios where the SD outperforms the MU method.
@@
Notation: Upper and lower boldface letters represent matrices and vectors, respectively. We define vector $\mathbf{v}=[\mathbf{v}_{1}^{T},\dots,\mathbf{v}_{K}^{T}]^{T}$, where $\mathbf{v}_{k}$ is the $k$-th block of vector $\mathbf{v}$ and $\mathbf{v}_{k,i}$ is the $i$-th element of the $k$-th block. $\mathbf{M}$ is a matrix comprising of submatrices, or blocks. $\mathbf{M}_{r,s,i,j}$ is the $(i,j)$-th element of the $r$-th block across and $s$-th block down. Matrices are indexed this way except where explicitly defined otherwise. $\mathbf{M}^{(k)}$ denotes a matrix relating to user $k$. $\mathbb{C}$ is the set of complex numbers. $\Re[\cdot]$ is the real operator. $\mathcal{CN}(\boldsymbol{\mu},\mathbf{Q})$ represents a complex Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance matrix $\mathbf{Q}$. $\mathbb{E}[\cdot]$ represents statistical expectation. $(\cdot)^{T}$, $(\cdot)^{*}$ and $(\cdot)^{\dagger}$ are the transpose, conjugate and Hermitian transpose operators, respectively. The angle of a vector, $\mathbf{x}$, of length $N$ is denoted as $\angle\mathbf{x}=[\angle{x}_{1},\dots,\angle{x}_{N}]^{T}$ and the exponent as $e^{\mathbf{x}}=[e^{{x}_{1}},\dots,e^{{x}_{N}}]^{T}$. $||\cdot||$ is the Euclidean norm. $_{1}F_{1}(a,b;z)$ is the confluent hypergeometric function and $_{2}F_{1}(a,b,c;z)$ is the Gaussian hypergeometric function. $L_{\nu}(\cdot)$ denotes a Laguerre function of non-integer degree $\nu$. $\Gamma(\cdot)$ is the complete gamma function. $f^{*}(r,\theta)$ is the probability density function (PDF) of a complex random variable $X=r\,\mathrm{e}^{j\theta}$, and $f^{*}(r_{1},r_{2},\theta_{1},\theta_{2})$ is the joint PDF of complex random variables $X_{1}=r_{1}\mathrm{e}^{j\theta_{1}}$ and $X_{2}=r_{2}\mathrm{e}^{j\theta_{2}}$. $J_{n}(\cdot)$ is the $n$-th order Bessel function of the first kind and $I_{n}(\cdot)$ is the $n$-th order modified Bessel function of the first kind.
@@ II System Model
We consider the UL RIS-aided system shown in Fig. 1, consisting of a RIS panel with $N$ elements, a BS with $M$ antennas, and $K$ single-antenna users located in the vicinity of the BS and RIS. A RIS link (UE-RIS-BS) and a direct link (UE-BS) connect the UE to the BS. The system bandwidth, $B$, is split into $K$ bands of $\frac{B}{K}$ Hz, with one user per band. The phase response of the RIS, and therefore the phase shift in the reflected signal, is fixed across the frequency bands. The RIS elements are grouped into “subsurfaces” and each subsurface is designed for a different user.
Figure 1: System model showing UL channels for UE 1 in band 1.
For UE $k$ in band $k$, let $\mathbf{h}^{(k)}_{\mathrm{d}}\!\in\mathbb{C}^{M\times 1}$, $\mathbf{h}^{(k)}_{\mathrm{ur}}\in\mathbb{C}^{N\times 1}$ and $\mathbf{H}^{(k)}_{\mathrm{rb}}\in\mathbb{C}^{M\times N}$ be the direct, UE-RIS and RIS-BS channels, respectively. $\mathbf{\Phi}\in\mathbb{C}^{N\times N}$ is a diagonal matrix of reflection coefficients for the RIS which can be given in block diagonal form, where the $k$-th block is designed to enhance the channel for user $k$. $\mathbf{\Phi}=\mathrm{diag}\{\mathbf{\Phi}_{1},\dots,\mathbf{\Phi}_{K}\}$, where $\mathbf{\Phi}_{k}=\mathrm{diag}(e^{j\phi_{k,1}},\dots,e^{j\phi_{k,N_{k}}})$. $N_{k}=N/K$ is the number of RIS elements chosen to support UE $k$ and $N=\sum^{K}_{k=1}N_{k}$. We consider an equal number of elements per UE, and non-equally divided subsurfaces are a potential topic for future work. Elements serving one UE are co-located, as this was demonstrated to be the best configuration in [16]. The received signal in band $k$ at the BS is
@@
where $s_{k}$ is the signal being sent from UE $k$, $\mathbb{E}[|s_{k}|^{2}]=E_{s}$, and $\mathbf{n}_{k}\sim\mathbb{C}\mathcal{N}(0,\sigma^{2}\mathbf{I})$ is additive white Gaussian noise (AWGN). Using the block diagonal form of $\mathbf{\Phi}$, (1) becomes
@@ II-A Channel Model
In this work, we consider correlated Ricean channels consisting of a rank-1 LoS and correlated Rayleigh component for all links. The LoS components are comprised of steering vectors and the scattered components adopt the Kronecker correlation model. Therefore, the channels for UE $k$ are
@@
where $\beta^{(k)}_{\mathrm{d}}$, $\beta^{(k)}_{\mathrm{rb}}$ and $\beta^{(k)}_{\mathrm{ur}}$ are the channel gains, $\kappa^{(k)}_{\mathrm{d}}$, $\kappa^{(k)}_{\mathrm{rb}}$ and $\kappa^{(k)}_{\mathrm{ur}}$ are the Ricean K-factors, $\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{LoS})}$, $\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{LoS})}$ and $\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{LoS})}$ are the LoS channel components for the $k$-th UE and $\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{SC})}$, $\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{SC})}$ and $\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{SC})}$ are the scattered channel (SC) components for the $k$-th UE. Here,
@@
This work focuses on strongly LoS $\mathbf{H}_{\mathrm{rb}}$ channels, where one ray is dominant but a scattered component is also present. This is a common and reasonable assumption for RIS systems, as the RIS can be located to exploit the LoS channel as seen in [19, 20, 21]. We also assume that reliable CSI is available for each individual link. Prior work has demonstrated that the UE-RIS and RIS-BS channels can be estimated separately, as in [22, 23]. Hence, we anticipate that analogous methods could be applied here.